OpenStudy (anonymous):

Assume the line of best fit for a data set is given by y=22x+32. What data points are predicted by the line of best fit? Check all that apply. a.(10,480) b.(5,142) c.(8,208) d.(7,154)

OpenStudy (mathteacher1729):

The idea here is to take each value of x and see if it gives the corresponding value of y. Example: Is (10,480) on the line y=22x+32 ? if YES then x = 10 y (should equal) 480 but... y = 22*(10) + 32 = 220 + 32 = 252 so x = 10 and y = 252 Therefore (10, 480) is NOT on the line because 252 is not 480. Does this help?

OpenStudy (anonymous):

yes. it sure does. thanks friend :)

OpenStudy (mathteacher1729):

Cool, glad it helped! :) If you use geogebra, you can very quickly see which points are and are not on the line. Make sure you can solve this problem both ways (graphing and doing algebra by hand). If you can, you've mastered it. :)

OpenStudy (anonymous):

thanks again C: